OpenStudy (anonymous):

An original rectangle has a length of ten and a width of twenty-eight. What happens to the area if the new width is one quarter the original width? Verify your answer.

jimthompson5910 (jim_thompson5910):

What is the area of the original rectangle?

OpenStudy (anonymous):

A = LW A' = L (W/4) = LW/4 =(10)(28)/4